OpenStudy (anonymous):

Solve for x: |x + 2| + 16 = 14

OpenStudy (mathmale):

Let's review "absolute value." What do you know about the absolute value function?

rebeccaxhawaii (rebeccaxhawaii):

^^

rebeccaxhawaii (rebeccaxhawaii):

absolute value is || signs ? what do they do?

OpenStudy (anonymous):

@mathmale Not much. I know that if an equation is equal to a positive number you'll get two solutions and things like that. I don't know how to begin solving this.

OpenStudy (mathmale):

Another way in which to ask my question: Are there ay restricions on the output of absolutel value expressions?

OpenStudy (mathmale):

Domain: all real numbers. Range: ... ?

OpenStudy (mathmale):

How would you respond if I gave you this: {x-1}=-5 ??

OpenStudy (anonymous):

@mathmale I would say that I don't know how to do this. I know nothing about math, honestly...

OpenStudy (mathmale):

The abs. value function MUST return a result that is either zero or greater than zero. So: (x-1)=-5 is false / impossible. No solutions. You are given |x + 2| + 16 = 14

OpenStudy (mathmale):

We'll try to isolate x. to do this, subtract 16 from both sides. What result do you get?

OpenStudy (anonymous):

-14?

OpenStudy (mathmale):

Yes. Think carefully. Is that OK?

OpenStudy (mathmale):

Review what I said about the absolute value function earlier.

OpenStudy (anonymous):

@mathmale thanks for the help! I got one of my teachers to explain it though and I understand it better now. (: